You are trying to swim as many laps as your friend did. After 10 minutes, you have 25 laps left. After 15 minutes, you have 15 laps left. How many laps did your friend swim? How much longer will it take you?
Answer by ankor@dixie-net.com(22740)

You are trying to swim as many laps as your friend did.
After 10 minutes, you have 25 laps left.
After 15 minutes, you have 15 laps left.
How many laps did your friend swim?
How much longer will it take you?
:
let x = the number of lap your friend swam
:
10 min 25 laps left
15 min 15 laps left
It took you 5 min to swim 10 laps, therefore 2 laps/min is your speed
:
x - 25 = 10(2)
x - 25 = 20
x = 20 + 25
x = 45 laps swam by your friend
;
your total time to swim 45 laps
45/2 = 22.5 min
:
How much longer will it take you? Assume this means after the 15 min point
22.5 - 15 = 7.5 min to finish